Circular No. 6321

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

GRO J1744–28 P. Blanco, University of California, San Diego; C. Lidman, European

Southern Observatory (ESO); and K. Glazebrook, Anglo-Australian Observatory, report photometry in the K band (2.2 μm) for objects in the vicinity of the radio source candidate for GRO J1744–28 (IAUC 6307, 6308): “These measurements were made on a co-added image (600 s total exposure) taken on Feb. 8 UT at the ESO 2.2-m telescope (+ IRAC2); this image provided the preliminary K-band flux and source identifications reported on IAUC 6315. Using the same nomenclature, we have determined the following K magnitudes: star A, 13.1; star B, 14.0; star C, 13.3; star D, 11.3. At this wavelength, 0.0 mag corresponds to 620.0 Jy. To aid calibration in the ongoing searches for near-infrared variability, we also report magnitudes for three nearby isolated stars in our image: α = 17^{h}44^{m}34^{s}.0, δ = −28^{o}45′07′′ (equinox 2000.0), K = 11.86; position end figures 37^{s}.3, 50′′, 11.75; 37^{s}.6, 12′′, 11.88. Although our absolute fluxes have uncertainties of ±0.25 mag, photometry of sources A–D relative to these brighter stars should be accurate to ±0.1 mag, except for star B (±0.3 mag), which is faint and confused with star D.”

G. V. Jung and B. F. Phlips, Universities Space Research Association; J. D. Kurfess, J. E. Grove, M. S. Strickman, W. N. Johnson, C. D. Dermer, and G. H. Share, Naval Research Laboratory (NRL); and S. J. Sturner, NRL and National Research Council, on behalf of the Gamma Ray Observatory OSSE team, report: “OSSE has observed the 467-ms pulsed emission from the bursting x-ray pulsar GRO J1744–28 (IAUC 6285), in both the steady source and during bursts. Using the pulsar ephemeris supplied by BATSE (M. Finger, private communication), we epoch-folded and co-added data from the brightest 2 s of each burst in a sample of 104 bursts observed by OSSE between Jan. 18 and 30. We find that the pulsed emission during the bursts lags the non-bursting pulsed emission; the lag amounts to 0.155 ± 0.007 of a pulsar period at the peak of the burst. The magnitude of the phase lag follows the average intensity profile of the burst sample. This result suggests that the burst-emitting region is physically offset from the region that produces the persistent emission, which could be caused by azimuthal asymmetries from accreting matter constrained to move along different field lines.”

1996 February 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green